OpenStudy (anonymous):

Find the equation of the inverse f(x)=4x-1

OpenStudy (anonymous):

f-1(x) = (x+1)/4

OpenStudy (anonymous):

it would be y =4x-1 then inverse is x=4y-1

OpenStudy (anonymous):

let y=4x-1 make x subject of formula (y+1)/4=x replace the y by x therefore f inverse of x=(x+1)/4